# Point2RBox-v3: Self-Bootstrapping from Point Annotations via Integrated Pseudo-Label Refinement and Utilization

### v2

@@ 1 INTRODUCTION
Oriented object detection (OOD) has attracted growing attention due to the increasing demand for object direction estimation in diverse fields, including autonomous driving (Feng et al., 2021), aerial images (Fu et al., 2020; Liu et al., 2017; Xia et al., 2018; Yang & Yan, 2022; Yang et al., 2018), scene text (Liao et al., 2018; Liu et al., 2018; Zhou et al., 2017), retail (Goldman et al., 2019; Pan et al., 2020), and industrial inspection (Liu et al., 2020; Wu et al., 2022).
@@
Training OOD models requires annotations as supervisory signals. Traditional rotated bounding boxes (RBoxes) provide accurate supervision but are costly: annotating each RBox is 36.5% more expensive than a horizontal box (HBox) and 104.8% more than a point (Yu et al., 2024), highlighting the potential of weakly-supervised OOD. HBox-based approaches such as H2RBox (Yang et al., 2023) and H2RBox-v2 (Yu et al., 2023; 2025b) have shown strong results, while point-based methods are rapidly advancing.
@@
Current point-supervised methods fall into four categories: (1) Pseudo generation using multiple instance learning and class probability maps (Luo et al., 2024; Ren et al., 2024); (2) Knowledge combination with one-shot learning (Yu et al., 2024); (3) Point-prompt OOD leveraging SAM’s zero-shot capability (Cao et al., 2023; Zhang et al., 2024a; Liu et al., 2025b; Lu & Bie, 2025; Kirillov et al., 2023); (4) Spatial layout for pseudo label generation (Yu et al., 2025a). Despite these advances, pseudo label quality remains a bottleneck, which motivates our approach.
@@
Motivation. Our motivation mainly stems from the lack of quality and utilization efficiency of pseudo labels in all current end-to-end methods. End-to-end point supervision method requires label assignment for the Feature Pyramid Network (FPN) (Lin et al., 2017a), which plays a crucial role in terms of detecting scale information. Most existing methods simply distribute them to the same layer, such as Point2RBox-v2 (Yu et al., 2025a). This wastes the scale information contained in the pseudo labels generated by the model. Point2RBox-v2 benefits from the “Voronoi Watershed Loss” for generating masks as pseudo labels. However, the watershed algorithm is less effective in sparse scenes. As SAM can also provide masks, we investigated it as an alternative but found that SAM fails in dense scenes. Can we enhance the detection capability of the model by focusing on these two aspects? In this paper, we have an in-depth discussion about this issue to explore how to fully uncover the potential of pseudo labels.
@@
Highlights. 1) Point2RBox-v3 is proposed for point-supervised OOD, advancing the state of the art (SOTA) as displayed in Figure 1 and Tables 4.1-4.2. 2) We improve the quality and utilization efficiency of pseudo labels, compared to Point2RBox-v2, without incurring a significant increase in training costs, which reveals more possibilities for the application and quality of pseudo labels in weakly-supervised object detection for future researchers.
@@
Contributions. 1) To our knowledge, Point2RBox-v3 is the first end-to-end point supervision model to explore how to assign pseudo labels to multiple FPN levels, and it complements the advantages of the SAM model with the watershed algorithm, which achieves excellent performance in both sparse and dense scenes. 2) We extend our method to partial weakly-supervised tasks (Liu et al., 2025a) beyond point supervision, demonstrating its adaptability and scalability. 3) The training pipeline and detailed implementation are elucidated. The source code will be made publicly available.
@@ 2 Related Work
Figure 2: The training pipeline of Point2RBox-v3. Progressive Label Assign utilizes scale info from pseudo label to dynamically assign gt point (see Figure 3). Prior-Guided Dynamic Mask provides enhanced mask supervision information (see Figure 4). $\mathcal{L}_{others}$ are the loss functions inherited from Point2RBox-v2 (see Appendix A.1).
Point-supervised Oriented Detection. Due to the low-cost and high-quality requirements in the field of object detection, point-supervised oriented object detection has become one of the research focuses. Some models apply the powerful zero-shot segmentation performance of the Segment Anything Model (SAM) (Kirillov et al., 2023) in their pipelines, such as P2RBox (Cao et al., 2023), PMHO (Zhang et al., 2024b), PointSAM (Liu et al., 2024) and PMS-SAM-RSD (Lu & Bie, 2025). Apart from these models, there are also some other models whose strategies involve creating pseudo labels using traditional machine learning or other methods. For example, the PointOBB series (Luo et al., 2024; Ren et al., 2025; Zhang et al., 2025b) employs techniques such as multi-instance learning and class probability graphs to generate pseudo RBoxes. Differently, Point2RBox (Yu et al., 2024) acquires knowledge from single sample examples through a knowledge combination method, and Point2RBox-v2 (Yu et al., 2025a) applies traditional machine learning algorithms such as watershed and edge detection. The technology based on SAM can achieve simpler applications by leveraging powerful base models, while other methods offer a solution that does not require reliance on other components. However, existing approaches have not yet explored their combination to enhance model performance.
Label Assignment. Point-supervised oriented object detection lacks scale information, and therefore cannot directly apply label assignment. PointSAM (Liu et al., 2025b) adopts a ViT (Dosovitskiy et al., 2020) backbone. Point2RBox (Yu et al., 2024), Point2RBox-v2 (Yu et al., 2025a) adopt only a single FPN feature level. PointOBB (Zhang et al., 2025a) selects grid points within the central region around each ground-truth point across all feature levels as positive samples. PointOBB-v3 (Zhang et al., 2025b) applies a gating mechanism to each FPN feature layer during training to produce corresponding gating scores, which are then used to automatically aggregate the multiple FPN output layers into a fused feature map for feature extraction. Unfortunately, these methods discard the classical multi-level label assignment in FPNs. We show that this omission largely explains the gap between point and full supervision, and that leveraging coarse scale cues from pseudo labels can effectively narrow this gap.
@@ 3.1 Overview and Preliminary
The architectural overview of Point2RBox-v3 is presented in Figure 2. Point2RBox-v3 retains the foundational components of Point2RBox-v2 (Yu et al., 2025a), which are ResNet50 (He et al., 2016) backbone, FPN (Lin et al., 2017a) head, PSC (Yu & Da, 2023) angle coder as well as losses ($\mathcal{L}_{others}$) shown in the figure. The details of $\mathcal{L}_{others}$ are provided in Appendix A.1. Our contribution lies in two enhanced modules: Progressive Label Assignment (PLA) and Prior-Guided Dynamic Mask Loss (PGDM-Loss).
@@ 3.2 Progressive Label Assignment
For both anchor-free (Tian et al., 2019) and anchor-based (Lin et al., 2017b) object detection algorithms, the scale information of objects plays an indispensable role in label assignment. Points lack scale, rendering label assignment infeasible. The scarcity of scale information persists and is even magnified in the loss constraints for scale learning. In Point2RBox-v2, the authors employ various approaches for scale learning, including explicit methods such as copy-paste loss, Voronoi Watershed loss, and edge loss, as well as implicit techniques like layout loss. Among these, watershed segmentation algorithm plays the most pivotal role in the scale learning task of the entire framework. Inspired by this, we introduce the pseudo labels originally used for scale constraints in the loss function into the label assignment module to provide approximate scale information.
@@
where Voronoi (Aurenhammer, 1991) is a partitioning of a space based on a set of points; $X$ are the annotated points within a training image; $V$ are the output ridges; $S$ are the output basin regions (pixel coordinates) corresponding to each annotated instance; $PL$ is for Pseudo Label.
@@
where $PL_{g}$ is the Pseudo Label for ground truth point $g$; $C_{g}$ refers to the candidate prediction boxes, selected by picking the prediction box associated with the anchor point closest to the ground-truth point $g$ at each FPN level. This approach can partially mitigate label assignment issues caused by solely relying on fixed segmentation regions generated by the watershed algorithm.
The improvement of the above design is verified in the Table 5 of the ablation study. Algorithm LABEL:alg:algorithm-pla in Appendix A.2 outlines the workings of the proposed method. Figure 3 illustrates the limitations of using fixed segmentation regions, such as those generated by watershed, for label assignment scale information. It also shows how, in the later stages of training, the dynamic forward predictions of the network progressively improve label assignment accuracy.
@@ 3.3 Prior-Guided Dynamic Mask Loss
Point2RBox-v2 (Yu et al., 2025a) leverages a Watershed Loss to generate pseudo labels from spatial layouts. While effective in dense scenes, this approach struggles in sparse scenarios where spatial cues are minimal, often leading to over- or under-segmentation. A detailed analysis of this limitation is provided in Appendix A.3. Conversely, the Segment Anything Model (SAM) (Kirillov et al., 2023) offers greater robustness in these challenging scenarios, but its high computational cost and weaker performance in dense scenes (Cai et al., 2024) preclude its direct application.
To harness the strengths of both methods, we propose the PGDM-Loss ($\mathcal{L}_{PGDM}$), a hybrid loss that dynamically routes images for mask generation. Images with a sparse instance count (total instances $\leq N_{thr}$) are directed to a SAM branch, while denser scenes are processed by the original -Watershed branch (see Fig. 4). This strategy enhances segmentation accuracy for difficult cases without a significant computational overhead, preserving the efficiency of the Voronoi Watershed method in dense scenes. We choose the lightweight MobileSAM (Zhang et al., 2023) as our SAM model. On DOTA-v1.0, its AP50 is the same as the result obtained using the basic SAM model, but its training time is 10 hours less than that of the basic one.
We integrate SAM not as a simple mask proposal tool, but as a source of weak supervision. Therefore, the SAM model does not participate in the inference process at all, ensuring its fast inference speed. For an instance $j$ of class $c_{j}$ routed to the SAM branch, SAM generates a set of candidate masks, $M_{j}=\{m_{1},m_{2},\dots,m_{k}\}$. We select the optimal mask $m_{j}^{*}$ via a prior-guided class-aware filtering mechanism governed by a scoring function:
@@
where $\phi_{k}(m_{i})$ represents metrics calculated from the masks, and $w_{k,c_{j}}$ is a class-specific weight based on simple prior knowledge (e.g., expected shape). A full description of the five metrics (center alignment, color consistency, rectangularity, circularity, and aspect ratio reliability) is available in Appendix A.4.
After obtaining masks from both branches, we compute the losses uniformly, following Point2RBox-v2 (Yu et al., 2025a). First, the mask $S$ is rotated to align with the current prediction to obtain the regression targets $w_{t}$ and $h_{t}$:
@@
The width-height regression loss for a single instance, which we denote as $\mathcal{L}_{mask}$, is then computed using the Gaussian Wasserstein Distance loss ($\mathcal{L}_{GWD}$) (Yang et al., 2021):
@@
where each $\mathcal{L}_{j}$ is computed for the corresponding instance using its optimal mask according to the formula for $\mathcal{L}_{mask}$ in Eq. equation 7.
@@ 4 Experiments
All experiments were conducted using PyTorch (Paszke et al., 2019) and MMRotate (Zhou et al., 2022). ResNet50 (He et al., 2016) was used as the backbone, and the model was trained with the AdamW (Loshchilov & Hutter, 2018) optimizer for 12 epochs. The initial learning rate was set to $5\times 10^{-5}$. We report the Average Precision (AP) on several remote sensing datasets, including DOTA (Xia et al., 2018), DIOR (Cheng et al., 2022), STAR (Li et al., 2024), and RSAR (Zhang et al., 2025c), without using multi-scale techniques. The Appendix A.5 contains detailed information on hyperparameters, data processing, and datasets.
@@
RepPoints (2019) 86.7 81.1 41.6 62.0 76.2 56.3 75.7 90.7 80.8 85.3 63.3 66.6 59.1 67.6 33.7 68.45
RetinaNet (2017b) 88.2 77.0 45.0 69.4 71.5 59.0 74.5 90.8 84.9 79.3 57.3 64.7 62.7 66.5 39.6 68.69
GWD (2021) 89.3 75.4 47.8 61.9 79.5 73.8 86.1 90.9 84.5 79.4 55.9 59.7 63.2 71.0 45.4 71.66
FCOS (2019) 89.1 76.9 50.1 63.2 79.8 79.8 87.1 90.4 80.8 84.6 59.7 66.3 65.8 71.3 41.7 72.44
S2A-Net (2022) 89.2 83.0 52.5 74.6 78.8 79.2 87.5 90.9 84.9 84.8 61.9 68.0 70.7 71.4 59.8 75.81
@@
Sun et al. (2021) $\times$ 51.5 38.7 16.1 36.8 29.8 19.2 23.4 83.9 50.6 80.0 18.9 50.2 25.6 28.7 25.5 38.60
BoxInst-RBox (2021)2 $\times$ 68.4 40.8 33.1 32.3 46.9 55.4 56.6 79.5 66.8 82.1 41.2 52.8 52.8 65.0 30.0 53.59
H2RBox (2023) 88.5 73.5 40.8 56.9 77.5 65.4 77.9 90.9 83.2 85.3 55.3 62.9 52.4 63.6 43.3 67.82
EIE-Det (2024) 87.7 70.2 41.5 60.5 80.7 76.3 86.3 90.9 82.6 84.7 53.1 64.5 58.1 70.4 43.8 70.10
H2RBox-v2 (2023) 89.0 74.4 50.0 60.5 79.8 75.3 86.9 90.9 85.1 85.0 59.2 63.2 65.2 70.5 49.7 72.31
@@
Point2Mask-RBox (2023)2 $\times$ 4.0 23.1 3.8 1.3 15.1 1.0 3.3 19.0 1.0 29.1 0.0 9.5 7.4 21.1 7.1 9.72
P2BNet+H2RBox (2023) $\times$ 24.7 35.9 7.1 27.9 3.3 12.1 17.5 17.5 0.8 34.0 6.3 49.6 11.6 27.2 18.8 19.63
P2BNet+H2RBox-v2 (2023) $\times$ 11.0 44.8 14.9 15.4 36.8 16.7 27.8 12.1 1.8 31.2 3.4 50.6 12.6 36.7 12.5 21.87
P2RBox (2024) $\times$ 87.8 65.7 15.0 60.7 73.0 71.7 78.9 81.5 44.5 81.2 41.2 39.3 45.5 57.5 41.2 59.04
PointOBB (2024) $\times$ 26.1 65.7 9.1 59.4 65.8 34.9 29.8 0.5 2.3 16.7 0.6 49.0 21.8 41.0 36.7 30.08
Point2RBox (2024) 62.9 64.3 14.4 35.0 28.2 38.9 33.3 25.2 2.2 44.5 3.4 48.1 25.9 45.0 22.6 34.07
Point2RBox+SK (2024) 53.3 63.9 3.7 50.9 40.0 39.2 45.7 76.7 10.5 56.1 5.4 49.5 24.2 51.2 33.8 40.27
Point2RBox+SK (2024) $\times$ 66.4 59.5 5.2 52.6 54.1 53.9 57.3 90.8 3.2 57.8 6.1 47.4 22.9 55.7 40.5 44.90
PointOBB-v2 (2025) $\times$ 64.5 27.8 1.9 36.2 58.8 47.2 53.4 90.5 62.2 45.3 12.1 41.7 8.1 43.7 32.0 41.68
PMS-SAM-RSD (2025) 69.0 39.5 6.7 44.8 64.7 71.9 79.6 79.8 2.7 60.0 12.1 32.6 39.6 44.8 42.5 46.00
PointOBB-v3 (2025b) 30.9 39.4 13.5 22.7 61.2 7.0 43.1 62.4 59.8 47.3 2.7 45.1 16.8 55.2 11.4 41.29
PointOBB-v3 (2025b) $\times$ 52.9 54.4 21.3 52.7 65.6 44.9 67.8 87.2 26.7 73.4 32.6 53.3 39.0 56.4 10.2 49.24
Point2RBox-v2 (2025a) 78.4 52.7 8.3 40.9 71.0 60.5 74.7 88.7 65.5 72.1 24.4 26.1 30.1 50.7 21.0 51.00
Point2RBox-v2 (2025a) $\times$ 88.0 72.6 8.0 46.2 79.6 76.3 86.9 89.1 79.7 82.9 26.2 45.3 45.8 66.3 46.3 62.61
[…]

### v3

@@ 1 INTRODUCTION
Oriented object detection (OOD) has attracted growing attention due to the increasing demand for object direction estimation in diverse fields, including autonomous driving [feng2021deep], aerial images [fu2020rotation, liu2017high, xia2018dota, yang2022arbitrary, yang2018automatic], scene text [liao2018rotation, liu2018fots, zhou2017east], retail [goldman2019precise, pan2020dynamic], and industrial inspection [liu2020data, wu2022pcbnet].
@@
Training OOD models requires annotations as supervisory signals. Traditional rotated bounding boxes (RBoxes) provide accurate supervision but are costly: annotating each RBox is 36.5% more expensive than a horizontal box (HBox) and 104.8% more than a point [yu2024point2rbox], highlighting the potential of weakly-supervised OOD. HBox-based approaches such as H2RBox [yang2023h2rbox] and H2RBox-v2 [yu2023h2rboxv2, yu2025wholly] have shown strong results, while point-based methods are rapidly advancing.
@@
Current point-supervised methods fall into four categories: (1) Pseudo generation using multiple instance learning and class probability maps [luo2024pointobb, pointobbv2]; (2) Knowledge combination with one-shot learning [yu2024point2rbox]; (3) Point-prompt OOD leveraging SAM’s zero-shot capability [cao2023p2rbox, 10649581, liu2025pointsam, lu2025semantic, kirillov2023segany]; (4) Spatial layout for pseudo label generation [yu2025point2rbox]. Despite these advances, pseudo label quality remains a bottleneck, which motivates our approach.
@@
Motivation. Our motivation mainly stems from the lack of quality and utilization efficiency of pseudo labels in all current end-to-end methods. End-to-end point supervision method requires label assignment for the Feature Pyramid Network (FPN) [lin2017feature], which plays a crucial role in terms of detecting scale information. Most existing methods simply distribute them to the same layer, such as Point2RBox-v2 [yu2025point2rbox]. This wastes the scale information contained in the pseudo labels generated by the model. Point2RBox-v2 benefits from the “Voronoi Watershed Loss” for generating masks as pseudo labels. However, the watershed algorithm is less effective in sparse scenes. As SAM can also provide masks, we investigated it as an alternative but found that SAM fails in dense scenes. Can we enhance the detection capability of the model by focusing on these two aspects? In this paper, we have an in-depth discussion about this issue to explore how to fully uncover the potential of pseudo labels.
@@
Highlights. 1) Point2RBox-v3 is proposed for point-supervised OOD, advancing the state of the art (SOTA) as displayed in Figure 1 and Tables 1-2. 2) We improve the quality and utilization efficiency of pseudo labels, compared to Point2RBox-v2, without incurring a significant increase in training costs, which reveals more possibilities for the application and quality of pseudo labels in weakly-supervised object detection for future researchers.
@@
Contributions. 1) To our knowledge, Point2RBox-v3 is the first end-to-end point supervision model to explore how to assign pseudo labels to multiple FPN levels, and it complements the advantages of the SAM model with the watershed algorithm, which achieves excellent performance in both sparse and dense scenes. 2) We extend our method to partial weakly-supervised tasks [liu2025partial] beyond point supervision, demonstrating its adaptability and scalability. 3) The training pipeline and detailed implementation are elucidated. The source code will be made publicly available.
@@ 2 Related Work
Figure 2: The training pipeline of Point2RBox-v3. Progressive Label Assign utilizes scale info from pseudo label to dynamically assign gt point (see Figure 3). Prior-Guided Dynamic Mask provides enhanced mask supervision information (see Figure 4). $\mathcal{L}_{others}$ are the loss functions inherited from Point2RBox-v2 (see Appendix 6.2).
Point-supervised Oriented Detection. Due to the low-cost and high-quality requirements in the field of object detection, point-supervised oriented object detection has become one of the research focuses. Some models apply the powerful zero-shot segmentation performance of the Segment Anything Model (SAM) [kirillov2023segany] in their pipelines, such as P2RBox [cao2023p2rbox], PMHO [zhang2024pmho], PointSAM [liu2024pointsam] and PMS-SAM-RSD [lu2025semantic]. Apart from these models, there are also some other models whose strategies involve creating pseudo labels using traditional machine learning or other methods. For example, the PointOBB series [luo2024pointobb, ren2024pointobbv2, zhang2025pointobbv3] employs techniques such as multi-instance learning and class probability graphs to generate pseudo RBoxes. Differently, Point2RBox [yu2024point2rbox] acquires knowledge from single sample examples through a knowledge combination method, and Point2RBox-v2 [yu2025point2rbox] applies traditional machine learning algorithms such as watershed and edge detection. The technology based on SAM can achieve simpler applications by leveraging powerful base models, while other methods offer a solution that does not require reliance on other components. However, existing approaches have not yet explored their combination to enhance model performance.
Label Assignment. Point-supervised oriented object detection lacks scale information, and therefore cannot directly apply label assignment. PointSAM [liu2025pointsam] adopts a ViT [dosovitskiy2020image] backbone. Point2RBox [yu2024point2rbox], Point2RBox-v2 [yu2025point2rbox] adopt only a single FPN feature level. PointOBB [zhang2025pointobb] selects grid points within the central region around each ground-truth point across all feature levels as positive samples. PointOBB-v3 [zhang2025pointobbv3] applies a gating mechanism to each FPN feature layer during training to produce corresponding gating scores, which are then used to automatically aggregate the multiple FPN output layers into a fused feature map for feature extraction. Unfortunately, these methods discard the classical multi-level label assignment in FPNs. We show that this omission largely explains the gap between point and full supervision, and that leveraging coarse scale cues from pseudo labels can effectively narrow this gap.
@@ 3.1 Overview and Preliminary
In brief, the definition of the Point-supervised Object Detection task is as follows: during the training phase, the model inputs consist of an image $I$ and the point annotations $P=\{(x_{i},y_{i})\}$ for all instances within the image (these points are typically, though not strictly, defined as the center points of instances). The training objective is to output the Rotated Bounding Box (RBox) representation $(x,y),(w,h),\theta]$ and the instance category $[cls]$ for each instance, where $(x,y),(w,h),\theta]$ denotes the precise circumscribed rectangle of the instance. During the inference phase, the input is solely the image $I$, and the model outputs the RBoxes $(x,y),(w,h),\theta]$ of all potential instances in the image, along with their corresponding categories $[cls]$ and confidence scores. The core difficulty of this task lies in overcoming the extreme deficiency of scale and orientation information inherent in point annotations. Specifically, the model is required to “reconstruct” precise geometric information from scratch. This requires resolving mutual boundary interference in densely packed scenes and addressing segmentation difficulties in sparse scenes caused by the scarcity of spatial constraints, while overcoming the challenge of diversity in scales.
The architectural overview of Point2RBox-v3 is presented in Figure 2. The simple evolutionary context of Point2RBox series is as follows: Point2RBox [yu2024point2rbox] established the angle prediction module via symmetry learning, while Point2RBox-v2 [yu2025point2rbox] introduced the scale prediction module based on spatial layout learning. The detailed explanation is in Appendix 6.1. While retaining the foundational components of this strong baseline (i.e., ResNet50 [he2016deep] backbone, FPN [lin2017feature] head, PSC [yu2023psc] angle coder, and losses $\mathcal{L}_{others}$ as shown in the Figure 2), Point2RBox-v3 focuses on two core innovations: Progressive Label Assignment (PLA) and Prior-Guided Dynamic Mask Loss (PGDM-Loss). These modules not only reinforce scale learning but, crucially, PLA enables the utilization of dynamic scale information for multi-level label assignment within the FPN under a weakly-supervised framework.
@@ 3.2 Progressive Label Assignment
For both anchor-free [tian2019fcos] and anchor-based [lin2017focal] object detection algorithms, the scale information of objects plays an indispensable role in label assignment. Points lack scale, rendering label assignment infeasible. The scarcity of scale information persists and is even magnified in the loss constraints for scale learning. In Point2RBox-v2, the authors employ various approaches for scale learning, including explicit methods such as copy-paste loss, Voronoi Watershed loss, and edge loss, as well as implicit techniques like layout loss. Among these, watershed segmentation algorithm plays the most pivotal role in the scale learning task of the entire framework. Inspired by this, we introduce the pseudo labels originally used for scale constraints in the loss function into the label assignment module to provide approximate scale information.
@@
where Voronoi [aurenhammer1991voronoi] is a partitioning of a space based on a set of points; $X$ are the annotated points within a training image; $I$ is the input image; $V$ are the output ridges; $S$ are the output basin regions (pixel coordinates) corresponding to each annotated instance; minAreaRect is used to calculate the minimum enclosing rotated rectangle of a point set; $PL$ is for Pseudo Label.
@@
where $PL_{g}$ is the Pseudo Label for ground truth point $g$; $C_{g}$ refers to the candidate prediction boxes, selected by picking the prediction box associated with the anchor point closest to the ground-truth point $g$ at each FPN level; function $score(b)$ is the raw classification confidence directly output by the detection network. This approach can partially mitigate label assignment issues caused by solely relying on fixed segmentation regions generated by the watershed algorithm.
The improvement of the above design is verified in the Table 5 of the ablation study. Algorithm 1 in Appendix 6.3 outlines the workings of the proposed method. Figure 3 illustrates the limitations of using fixed segmentation regions, such as those generated by watershed, for label assignment scale information. It also shows how, in the later stages of training, the dynamic forward predictions of the network progressively improve label assignment accuracy.
@@ 3.3 Prior-Guided Dynamic Mask Loss
Point2RBox-v2 [yu2025point2rbox] leverages a Watershed Loss to generate pseudo labels from spatial layouts. While effective in dense scenes, this approach struggles in sparse scenarios where spatial cues are minimal, often leading to over- or under-segmentation. A detailed analysis of this limitation is provided in Appendix 6.4. Conversely, the Segment Anything Model (SAM) [kirillov2023segany] offers greater robustness in these challenging scenarios, but its high computational cost and weaker performance in dense scenes [cai2024crowd] preclude its direct application.
To harness the strengths of both methods, we propose the PGDM-Loss ($\mathcal{L}_{PGDM}$), a hybrid loss that dynamically routes images for mask generation. Images with a sparse instance count (total instances $\leq N_{thr}$) are directed to a SAM branch, while denser scenes are processed by the original -Watershed branch (see Figure 4). This strategy enhances segmentation accuracy for difficult cases without a significant computational overhead, preserving the efficiency of the Voronoi Watershed method in dense scenes. We choose the lightweight MobileSAM [mobile_sam] as our SAM model. On DOTA-v1.0, its AP50 is the same as the result obtained using the basic SAM model, but its training time is 10 hours less than that of the basic one.
We integrate SAM not as a simple mask proposal tool, but as a source of weak supervision. Therefore, the SAM model does not participate in the inference process at all, ensuring its fast inference speed. For an instance $j$ of class $c_{j}$ routed to the SAM branch, SAM generates a set of candidate masks, $M_{j}=\{m_{1},m_{2},\dots,m_{k}\}$. We select the optimal mask $m_{j}^{*}$ via a prior-guided filtering mechanism governed by a scoring function:
@@
where $\phi_{k}(m_{i})$ represents metrics calculated from the masks, and $w_{k,c_{j}}$ is a class-specific weight based on simple prior knowledge (e.g., expected shape). A full description of the five metrics (center alignment, color consistency, rectangularity, circularity, and aspect ratio reliability) is available in Appendix 6.5.
After obtaining masks from both branches, we compute the losses uniformly, following Point2RBox-v2 [yu2025point2rbox]. First, the mask $S$ is rotated to align with the current prediction to obtain the regression targets $w_{t}$ and $h_{t}$:
@@
The width-height regression loss for a single instance, which we denote as $\mathcal{L}_{mask}$, is then computed using the Gaussian Wasserstein Distance loss ($\mathcal{L}_{GWD}$) [yang2021rethinking]:
@@
where each $\mathcal{L}_{j}$ is computed for the corresponding instance using its optimal mask according to the formula for $\mathcal{L}_{mask}$ in Eq. 7.
@@ 4 Experiments
4.1 Experimental Details
Implementation Details. Experiments are carried out using PyTorch 1.13.1 [paszke2019pytorch] and the rotation detection toolkit MMRotate 1.0.0 [zhou2022mmrotate]. All experiments follow the same hyper-parameters. We adopt Average Precision (AP) as the primary evaluation metric. All models are based on the ResNet50 [he2016deep] backbone and are trained using AdamW [loshchilov2018decoupled].
  1. 1.
  2. 2.
    Epochs: 12 for all datasets.
  3. 3.
    Augmentation: Random flip for all datasets.
  4. 4.
  5. 5.
    Multi-scale: All experiments are evaluated without multi-scale techniques [zhou2022mmrotate].
  • DOTA [xia2018dota]. DOTA-v1.0 contains 2,806 aerial images with 15 categories. DOTA-v1.5/2.0 are its extended versions with more small objects and new categories.
  • DIOR [li2020object]. It is an aerial image dataset re-annotated with RBoxes based on its original HBox version [cheng2022anchor], featuring high variation in object scale and high intra-class diversity.
  • STAR [li2024star]. An extensive dataset for scene graph generation, covering over 210,000 objects with diverse spatial resolutions and 48 fine-grained categories.
  • RSAR [zhang2025rsar]. A remote sensing dataset based on Synthetic Aperture Radar (SAR) imagery, containing 6 categories.
@@ 4.1 Experimental Details
RepPoints [yang2019reppoints] 86.7 81.1 41.6 62.0 76.2 56.3 75.7 90.7 80.8 85.3 63.3 66.6 59.1 67.6 33.7 68.45
RetinaNet [lin2017focal] 88.2 77.0 45.0 69.4 71.5 59.0 74.5 90.8 84.9 79.3 57.3 64.7 62.7 66.5 39.6 68.69
GWD [yang2021rethinking] 89.3 75.4 47.8 61.9 79.5 73.8 86.1 90.9 84.5 79.4 55.9 59.7 63.2 71.0 45.4 71.66
FCOS [tian2019fcos] 89.1 76.9 50.1 63.2 79.8 79.8 87.1 90.4 80.8 84.6 59.7 66.3 65.8 71.3 41.7 72.44
S2A-Net [han2022align] 89.2 83.0 52.5 74.6 78.8 79.2 87.5 90.9 84.9 84.8 61.9 68.0 70.7 71.4 59.8 75.81
@@
Sun et al. [sun2021oriented] $\times$ 51.5 38.7 16.1 36.8 29.8 19.2 23.4 83.9 50.6 80.0 18.9 50.2 25.6 28.7 25.5 38.60
BoxInst-RBox [tian2021boxinst]2 $\times$ 68.4 40.8 33.1 32.3 46.9 55.4 56.6 79.5 66.8 82.1 41.2 52.8 52.8 65.0 30.0 53.59
H2RBox [yang2023h2rbox] 88.5 73.5 40.8 56.9 77.5 65.4 77.9 90.9 83.2 85.3 55.3 62.9 52.4 63.6 43.3 67.82
EIE-Det [wang2024explicit] 87.7 70.2 41.5 60.5 80.7 76.3 86.3 90.9 82.6 84.7 53.1 64.5 58.1 70.4 43.8 70.10
H2RBox-v2 [yu2023h2rboxv2] 89.0 74.4 50.0 60.5 79.8 75.3 86.9 90.9 85.1 85.0 59.2 63.2 65.2 70.5 49.7 72.31
@@
Point2Mask-RBox [li2023point2mask]2 $\times$ 4.0 23.1 3.8 1.3 15.1 1.0 3.3 19.0 1.0 29.1 0.0 9.5 7.4 21.1 7.1 9.72
P2BNet+H2RBox [yang2023h2rbox] $\times$ 24.7 35.9 7.1 27.9 3.3 12.1 17.5 17.5 0.8 34.0 6.3 49.6 11.6 27.2 18.8 19.63
P2BNet+H2RBox-v2 [yu2023h2rboxv2] $\times$ 11.0 44.8 14.9 15.4 36.8 16.7 27.8 12.1 1.8 31.2 3.4 50.6 12.6 36.7 12.5 21.87
P2RBox [yu2024point2rbox] $\times$ 87.8 65.7 15.0 60.7 73.0 71.7 78.9 81.5 44.5 81.2 41.2 39.3 45.5 57.5 41.2 59.04
PointOBB [luo2024pointobb] $\times$ 26.1 65.7 9.1 59.4 65.8 34.9 29.8 0.5 2.3 16.7 0.6 49.0 21.8 41.0 36.7 30.08
Point2RBox [yu2024point2rbox] 62.9 64.3 14.4 35.0 28.2 38.9 33.3 25.2 2.2 44.5 3.4 48.1 25.9 45.0 22.6 34.07
Point2RBox+SK [yu2024point2rbox] 53.3 63.9 3.7 50.9 40.0 39.2 45.7 76.7 10.5 56.1 5.4 49.5 24.2 51.2 33.8 40.27
Point2RBox+SK [yu2024point2rbox] $\times$ 66.4 59.5 5.2 52.6 54.1 53.9 57.3 90.8 3.2 57.8 6.1 47.4 22.9 55.7 40.5 44.90
PointOBB-v2 [ren2024pointobbv2] $\times$ 64.5 27.8 1.9 36.2 58.8 47.2 53.4 90.5 62.2 45.3 12.1 41.7 8.1 43.7 32.0 41.68
PMS-SAM-RSD [lu2025semantic] 69.0 39.5 6.7 44.8 64.7 71.9 79.6 79.8 2.7 60.0 12.1 32.6 39.6 44.8 42.5 46.00
PointOBB-v3 [zhang2025pointobbv3] 30.9 39.4 13.5 22.7 61.2 7.0 43.1 62.4 59.8 47.3 2.7 45.1 16.8 55.2 11.4 41.29
PointOBB-v3 [zhang2025pointobbv3] $\times$ 52.9 54.4 21.3 52.7 65.6 44.9 67.8 87.2 26.7 73.4 32.6 53.3 39.0 56.4 10.2 49.24
Point2RBox-v2 [yu2025point2rbox] 78.4 52.7 8.3 40.9 71.0 60.5 74.7 88.7 65.5 72.1 24.4 26.1 30.1 50.7 21.0 51.00
Point2RBox-v2 [yu2025point2rbox] $\times$ 88.0 72.6 8.0 46.2 79.6 76.3 86.9 89.1 79.7 82.9 26.2 45.3 45.8 66.3 46.3 62.61
Point2RBox-v2+PLA (ours) 82.3 47.8 23.2 36.0 77.9 75.5 86.6 85.3 72.0 76.5 28.0 33.4 39.4 52.7 31.7 56.55
Point2RBox-v2+PLA (ours) $\times$ 88.3 71.5 26.3 45.9 78.8 76.9 87.5 86.9 74.6 83.1 47.7 45.2 49.0 60.7 46.7 64.61
Point2RBox-v2+PLA+CS3 (ours) 84.6 51.5 25.4 37.2 78.7 75.3 86.5 86.4 75.8 74.9 31.3 33.6 40.2 54.5 34.2 58.00
Point2RBox-v3 (ours) 86.5 53.4 35.7 37.5 78.8 75.3 86.3 86.6 66.0 80.2 29.7 49.9 36.7 59.1 32.5 59.61
Point2RBox-v3 (ours) $\times$ 89.0 72.5 41.6 45.1 79.4 76.4 87.7 85.1 75.7 84.8 44.4 55.4 45.7 59.8 48.8 66.09